# Bimanual Regrasp Planning and Control for Eliminating Object Pose Uncertainty

### v1

@@ VI Experiments and Analysis
To ensure clarity in comparison and analysis, we used the simple L-shaped object shown in Fig. 7 and assumed a single initial pose in the planning simulation. The object has a long edge of $125$ mm and a short edge of $100$ mm, with a square cross-section of $25$ mm $\times$ $25$ mm. The surfaces of the L-shaped object are mutually perpendicular. They could align with the three axes of the coordinate system and facilitate analysis. The initial pose of the object in the simulation was set to $(700,300,1.25)$ mm. Its rotation was set to be aligned with the coordinate frame of the robot platform (recall the colored frame shown in Fig. 7). The long edge aligns with the $\boldsymbol{y}$ axis. The short edge aligns with the $\boldsymbol{z}$ axis. Similarly to the initial pose, we set the regrasp pose of the object in the simulation to be at a unique fixed location $(680,50.0,120.0)$ mm, with the rotation identical to the initial one to facilitate analysis. The location is in the middle of the two arms, making it easy to plan effective bimanual regrasp motions. The regrasp motion planned based on the parameter settings and the proposed planning methods led to orthogonal triples where the first selected grasp is along the $\boldsymbol{y}$ axis, the second grasp is along the $\boldsymbol{z}$ axis, and the third grasp is along the $x$ axis. Together, the three grasps uniquely constrain the object’s pose.
In the real world, we manually introduced uncertainties in multiple directions to the object for experimental evaluation. These uncertainties are illustrated in Fig. 8. Fig. 8(a) represents initial object poses obtained by introducing uncertainties in the horizontal directions (back, front, left, and right) relative to a center pose shown in Fig. 8(a.v). The center pose closely corresponds to the initial pose of the object used in the simulation. Note that the center pose is not exactly the initial pose in the simulation. The front, back, left, and right movements are approximate ±$\pm 12.5$ mm and are also not exact. We are not enforcing precise placement of the object when preparing these poses. The reason is that preparing exact uncertainty in the real world can be challenging. Finger contact during the grasping process naturally causes additional errors. Specifying exact uncertainty values and evaluating absolute accuracy would not provide meaningful insights. We analyze repeatability errors rather than absolute errors to evaluate the effectiveness of our method in eliminating uncertainties. In addition to the horizontal uncertainties, we also prepared initial poses with uncertainties in the upward, upper-front, and upper-back directions. These poses are shown in Fig. 8(b). The L-shape object is elevated using a white block while being shifted back and forward with the same approximate ±$\pm 12.5$ mm offset when preparing these uncertain poses.
Figure 8: (a) $9$ uncertain poses in the horizontal directions. (a.i) BL – Back Left. (a.ii) BC – Back Center. (a.iii) BR – Back Right. (a.iv) L – Left. (a.v) C – Center. (a.vi) R – Right. (a.vii) FL – Front Left. (a.viii) FC – Front Center. (a.ix) FR – Front Right. (b) $3$ uncertain poses in upward directions. (b.i) DC – Down Center (for reference). (b.ii) UC – Upper Center. (b.iii) Upper Back. (b.iv) Upper Front.
Using the planned robot motions, we performed regrasping and applied admittance control following the proposed method for each initial object pose in Figs. 8(a) and (b), and estimated their final object poses using equation 6. The estimation results are summarized in the “Proposed method” columns of Table I. The results corresponding to (a) and (b) are separated by a double solid line. From the results, we can observe that the estimation performs reasonably well. In the case of the uncertain object placement shown in Figs. 8(a), the estimated $\boldsymbol{p}_{o_{r}}$ vector has a mean and variance of $(9.82,0.002)$ mm and $(898.33,0.019)$ mm in the $\boldsymbol{y}$ and $\boldsymbol{z}$ directions respectively. The low variance aligns well with our intuition – For the uncertain poses listed in Fig. 8(a), the object is placed on the table. Its position in the $\boldsymbol{z}$-direction remains the same. Meanwhile, the object is always grasped by fingers closing from the $\boldsymbol{y}$-direction. Its uncertainty in the $\boldsymbol{y}$-direction is constrained first grasp’s two fingers. Therefore, the final estimated poses should be consistent (have a low variance) in these two directions. On the other hand, the first values of the estimated $\boldsymbol{p}_{o_{r}}$ can be clearly categorized into three distinct groups. Among them, BC and BR form one group, with a mean and variance of $(687.05,2.1)$ mm. L, C, and R constitute another group, with a mean and variance of $(699.3,1.4)$ mm. Finally, FL and FC belong to the third group, with a mean and variance of $(709.45,2.1)$ mm. This result is also consistent with our initial placement. The object displacement in the $\boldsymbol{x}$ direction can be clearly distinguished, and the repeatability is satisfying. The variance of each distinct group falls within the range of the manual placement uncertainty. Similar analysis and conclusions can be drawn from the results for the uncertain object placements shown in Figs. 8(b). They are in the UB, UC, and UF rows. The estimated $\boldsymbol{p}_{o_{r}}$ vectors have nearly the same values in the $\boldsymbol{y}$ and $\boldsymbol{z}$ directions and clearly distinct values in the $\boldsymbol{x}$ direction. For rotation, all estimated results are generally aligned with the coordinate axes. They have an average deviation of $0.59^{\circ}$ with $0.11^{\circ}$ variance from rotation matrix $\begin{bmatrix}\boldsymbol{x}&\boldsymbol{z}&-\boldsymbol{y}\end{bmatrix}$. This outcome is consistent with the fact that the object’s initial placement is always on a plane and that the opening-closing direction of the first grasp is along the $\boldsymbol{y}$ axis. Its repeatability is also satisfying considering the small variance. Note that the BL and FR rows in the table do not contain values. This is because the object’s pose uncertainty was too large, causing the robot to collide with the object during execution. The maximum tolerable uncertainty in our method is constrained by the gripper’s opening width. When the uncertainty exceeds half of the gripper’s opening width, collisions may occur and lead to execution failure. Additionally, the table does not include estimation results for DC pose. This pose is only provided to give readers a reference for comparison before the object is elevated, and thus, no estimation was performed for it.
TABLE I: Pose estimation results
Type Exec. Proposed method AR marker
$\boldsymbol{p}_{o_{r}}$ (mm) $\boldsymbol{R}_{o_{r}}$ $\boldsymbol{p}_{o_{r}}$ (mm) $\boldsymbol{R}_{o_{r}}$
BL Fail. - - - -
BC Succ. $(685.6,49.8,118.2)$ $\begin{bmatrix}1.00&-0.00&-0.00\\ -0.01&0.00&-1.00\\ -0.01&1.00&0.00\end{bmatrix}$ $(715.4,56.5,128.3)$ $\begin{bmatrix}1.00&-0.01&0.07\\ 0.07&-0.02&-1.00\\ 0.01&1.00&-0.02\end{bmatrix}$
BR Succ. $(688.5,49.8,118.3)$ $\begin{bmatrix}1.00&-0.00&-0.00\\ -0.02&-0.00&-1.00\\ -0.01&1.00&0.00\end{bmatrix}$ $(719.5,34.8,146.3)$ $\begin{bmatrix}1.00&-0.06&0.00\\ 0.00&-0.02&-1.00\\ 0.06&1.00&-0.02\end{bmatrix}$
L Succ. $(699.5,49.9,118.5)$ $\begin{bmatrix}1.00&-0.01&0.00\\ 0.00&-0.00&-1.00\\ -0.00&1.00&0.00\end{bmatrix}$ $(740.2,54.1,131.5)$ $\begin{bmatrix}1.00&-0.03&0.05\\ 0.05&-0.01&-1.00\\ 0.03&1.00&-0.00\end{bmatrix}$
C Succ. $(697.8,49.9,118.1)$ $\begin{bmatrix}1.00&-0.00&0.00\\ -0.00&-0.00&-1.00\\ -0.01&1.00&-0.00\end{bmatrix}$ $(726.9,36.8,144.5)$ $\begin{bmatrix}1.00&-0.07&-0.01\\ -0.01&-0.01&-1.00\\ 0.07&1.00&-0.01\end{bmatrix}$
R Succ. $(700.7,49.8,118.3)$ $\begin{bmatrix}1.00&-0.00&-0.00\\ -0.00&-0.00&-1.00\\ -0.01&1.00&0.00\end{bmatrix}$ $(726.3,55.6,145.8)$ $\begin{bmatrix}1.00&-0.07&0.07\\ 0.06&-0.01&-1.00\\ 0.07&1.00&-0.01\end{bmatrix}$
FL Succ. $(710.9,49.8,118.4)$ $\begin{bmatrix}1.00&0.01&-0.00\\ 0.03&-0.00&-1.00\\ -0.01&1.00&0.00\end{bmatrix}$ $(741.3,57.7,128.4)$ $\begin{bmatrix}0.99&-0.01&0.10\\ 0.10&-0.01&-0.99\\ 0.01&1.00&-0.01\end{bmatrix}$
FC Succ. $(708.0,49.8,118.5)$ $\begin{bmatrix}1.00&0.00&-0.00\\ 0.02&-0.00&-1.00\\ -0.01&1.00&-0.00\end{bmatrix}$ $(756.1,29.4,110.6)$ $\begin{bmatrix}1.00&-0.00&0.01\\ 0.01&-0.03&-1.00\\ 0.00&1.00&-0.03\end{bmatrix}$
FR Fail. - - - -
UB Succ. $(689.8,49.8,131.8)$ $\begin{bmatrix}1.00&0.01&0.00\\ -0.02&-0.01&-1.00\\ -0.03&1.00&0.00\end{bmatrix}$ $(717.5,38.3,157.9)$ $\begin{bmatrix}1.00&-0.06&-0.02\\ -0.02&-0.03&-1.00\\ 0.06&1.00&-0.03\end{bmatrix}$
UC Succ. $(698.7,49.8,131.6)$ $\begin{bmatrix}1.00&-0.01&0.00\\ 0.00&-0.00&-1.00\\ -0.00&1.00&0.00\end{bmatrix}$ $(732.4,32.1,160.6)$ $\begin{bmatrix}1.00&-0.07&0.02\\ 0.02&-0.02&-1.00\\ 0.07&1.00&-0.02\end{bmatrix}$
UF Succ. $(706.6,49.8,131.6)$ $\begin{bmatrix}1.00&-0.00&-0.00\\ 0.01&-0.02&-1.00\\ -0.01&1.00&0.00\end{bmatrix}$ $(735.6,36.6,158.5)$ $\begin{bmatrix}1.00&-0.08&0.02\\ 0.01&-0.02&-1.00\\ 0.08&1.00&-0.02\end{bmatrix}$
[…]
@@ VII Conclusions and Future Work
This paper proposes a method for eliminating object uncertainties by leveraging the grasping constraints of parallel grippers with flat finger pads and a bimanual regrasp planning method. We applied this method to estimate grasping errors for a roughly placed L-shaped object. The results demonstrate that the proposed approach effectively eliminates uncertainties and achieves high repeatability accuracy.
Currently, this method is only applicable when all three grasps involve planar contact. For curved surface contacts, additional grasps may be required to fully constrain uncertainties. Further exploring the necessary grasp combinations will be an important future topic.

### v3

@@ VI Experiments and Analysis
Fig. 8 shows the objects used in the experiments, their dimensions, and the randomly assigned placements in experimental trials. The objects are L-shaped items with square and diamond cross-sections, as described earlier. Their long arm measures 125 mm, the short arm 100 mm, and the distance between the parallel edges of the cross-section is 25 mm for both. The diamond-shaped cross-section has a 75 acute interior angle. Optical tracking markers were attached at pre-drilled holes on the objects. During grasp planning, we can automatically generate grasps and motions that avoid collisions with the markers. For each object, we randomly assigned three different initial placements, labeled L1$\sim$3 and TL1$\sim$3 in Fig. 8, and tasked the robots with moving the objects to predefined target positions. The supplementary video accompanying this paper contains details of the grasp selection, motion planning, and robot actions. We recommend readers to refer to this video for an intuitive understanding of the experimental procedures.
Figure 8: (a) L-shape object and its random initial placements. (b) TL-shape object and its random initial placements.
Table I summarizes the results. The data section has two main column groups reporting the positional differences and rotational differences. Within each group, there are three subcolumns. The $g_{2}$ columns show the positional and rotational differences of the $g_{2}$ grasp before and after admittance control. The $g_{3}$ columns present the corresponding differences for the $g_{3}$ grasp. The “object” columns report the differences between the object pose estimated by the proposed method and the pose measured by the optical tracking system. The results for $g_{1}$ are not included, as it involves no admittance control.
TABLE I: Experimental Results with Statistics
Task Position Differences ($\Delta\mathbf{p}$, mm) Rotation Vector Differences ($\Delta\mathbf{w}$, degrees)
$g_{2}$ $g_{3}$ object $g_{2}$ $g_{3}$ object
L1 -0.40 -1.60 5.20 -8.90 -2.00 -0.80 -33.70 -61.80 32.20 -0.55 -0.91 -0.02 -0.34 -0.75 -0.91 -0.01 2.82 -1.39
L2 -0.30 -1.70 5.20 -0.20 -2.00 -0.80 -33.60 -61.70 32.30 -0.51 -0.14 -0.03 -0.33 -0.25 -0.06 -0.12 1.83 -1.95
L3 -0.40 -1.60 5.20 -9.00 -1.90 -0.40 -33.20 -61.50 31.70 -0.30 -0.13 -0.00 -0.20 -0.51 -1.11 -0.00 0.98 -1.07
Mean -33.50 -61.67 32.07 -0.04 1.87 -1.47
Std -00.25 -00.15 00.25 -0.06 0.95 -0.44
TL1 -1.81 -0.99 0.38 -8.02 -8.05 -7.20 -59.35 -32.75 52.41 -0.64 -1.07 -1.00 -0.45 -0.36 -0.78 -0.01 2.05 -0.42
TL2 -0.08 -0.18 1.31 -0.40 -1.60 5.20 -59.66 -34.42 51.76 -0.00 -0.00 -0.00 -0.30 -0.13 -0.00 -0.13 0.73 -1.09
TL3 -0.07 -0.21 1.17 -1.71 -0.36 -1.47 -56.37 -36.34 52.26 -0.04 -0.01 -0.00 -1.25 -0.29 -0.56 -0.86 0.27 -2.11
Mean -58.46 -34.50 52.14 -0.33 1.02 -1.21
Std -01.86 -01.50 00.36 -0.47 0.38 -0.85
From the table, we observe that both the $g_{2}$ and $g_{3}$ grasps exhibit substantial changes before and after admittance control, indicating that the robot actively adjusts the regrasp pose to estimate uncertainties for each task. However, the final estimation results are not satisfactory, showing positional differences of up to 61.67 mm and rotational differences of up to 1.86 degrees compared to the optical tracking system.
Upon closer examination, we found that these errors are primarily systematic. For the L-shaped object, the standard deviations in $\Delta\mathbf{p}$ across tasks were [0.25, 0.15, 0.25] mm, all below 0.3 mm, and the deviations in $\Delta\mathbf{w}$ were [0.06, 0.95, 0.44] degrees, all below 1.0 degrees. Similarly, for the TL-shaped object, the deviations in $\Delta\mathbf{p}$ and $\Delta\mathbf{w}$ were [1.86, 1.50, 0.36] mm and [0.47, 0.38, 0.85] degrees, respectively, remaining below 2 mm and 0.9 degrees. These results suggest that while the overall estimation errors appear large, their internal consistency is high, indicating that the system’s inaccuracies stem largely from systematic biases. Likely sources include errors from the optical tracking system itself (due to the 3D-printed marker mounts for calibration and robot frame location), installation errors in the dual-arm robot, and absolute control errors of the robot. In addition, when comparing the deviations between the L-shaped and TL-shaped objects, we found that the inference results for the L-shaped object were better than those for the TL-shaped object. This outcome differs from our theoretical expectations. Theoretically, even if the three grasps are not perfectly orthogonal, the final inference accuracy should not degrade, and the L- and TL-shaped objects should exhibit similar deviations. We believe that the observed difference arises because the TL-shaped object involves rotations around axes unaligned with body frames, which place higher demands on the assembly precision of the dual-arm robot. The larger deviations reveal that the robot setup has greater misalignment in rotational components.
In conclusion, although the results exhibit large absolute errors, the deviation levels are on the same order of magnitude as the optical tracking system. The results indicate that the proposed method has strong relative pose estimation performance.
@@ VII Conclusions and Future Work
This paper proposes a method for reducing object uncertainties by leveraging the grasping constraints of parallel grippers with flat finger pads and a bimanual regrasp planning method. We applied this method to estimate grasping errors for a roughly placed L-shaped object. The results demonstrate that the proposed approach can effectively reduce uncertainties and achieves high repeatability accuracy.
The current method assumes planar contacts between flat fingertips and flat object surfaces. When the object surface is curved, the number of degrees of freedom constrained by a grasp will vary, and more grasps may be required to constrain uncertainties. Investigating the minimal and sufficient set of grasps under such conditions is an interesting future direction. Additionally, exploring learning-based methods to predict grasps and motion will be helpful to improve efficiency.